OpenStudy (anonymous):

\[V(x)=E(x^{2})-\mu^{2}\] then wat is V(-ax) in simplified form as i given for before.

OpenStudy (zarkon):

the -a gets pulled out and squared

OpenStudy (zarkon):

\[V(-aX)=(-a)^2V(X)=a^2V(X)\]

OpenStudy (zarkon):

it is easyier to see if write is as \[V(X)=E(X^2)-(E(X))^2\] and use the fact that the expected value is a linear opperator \[E(aX+b)=aE(X)+b\]